OpenStudy (anonymous):

Determine if the following Serie converges or not \[\Large{\sum_{k\in\mathbb{N}}\frac{(-1)^{k}}{k^{2}+1}}\]

OpenStudy (blockcolder):

I suggest the alternating series test for this one.

OpenStudy (anonymous):

how you should apply alternating series test to this serie ?

OpenStudy (anonymous):

*would

OpenStudy (blockcolder):

Let's say the series is in the form \(\sum (-1)^nb_n\). If \(b_n\) satisfies the following conditions: 1. \(b_n\) is decreasing, i.e. \(b_{n+1}<b_n\), and 2. \(\lim_{n \rightarrow \infty} b_n=0\), then the series \(\sum (-1)^n b_n\) converges.

OpenStudy (anonymous):

ist it Leipniz Criterium ?

OpenStudy (blockcolder):

Yeah, that's another name for this test.

OpenStudy (anonymous):

ok thx =)